# Complexity transitions in chaotic quantum systems

### v1

@@ V.2 Results
In the integrable Ising model at $\lambda=1$, the ground state is doubly degenerate. However, introducing a small perturbation from the SYK Hamiltonian, $\lambda\lesssim 1$, lifts this degeneracy. Due to the fermion parity symmetry of the full Hamiltonian, the system selects a specific ground state: an equal superposition of the two antiferromagnetic spin configurations, with a relative phase of ±$\pm 1$. As a result, in this limit, the fractal dimension and maximal von-Neumann entropy density approach the values of $1/N$. In contrast, the magic tends to zero since this superposition is a stabilizer state (see Appendix G). All considered markers approach the values of the $\lambda\sim 1$ limit relatively rapidly without appreciable changes beyond $\lambda\gtrsim 0.5$, see Fig. 6.
@@
where $\mathcal{C}$ denotes one of the three markers $D_{2},S_{\rm vN}$, or $\mathcal{M}_{2}$. As shown in the insets of Fig. 6, for sufficiently large $M$, $\langle\Delta D_{2}\rangle$ and $\langle\Delta S_{\rm vN}\rangle$ tend to zero as $\lambda\to 0$, whereas $\langle\Delta\mathcal{M}_{2}\rangle$ is strictly larger than zero. This behavior can be understood by the fact that, similarly to what we have seen for $\lambda\sim 1$, the ground state for small $\lambda$ has fixed fermion parity, which means that only Pauli strings composed of the product of an even number of Majorana operators (those that preserve the symmetry) contribute to a non-zero $\mathcal{M}_{2}$. Since these comprise only half of all possible Pauli strings, the overall $\mathcal{M}_{2}$ is smaller than the Haar-random value. The situation is, however, a bit more subtle in the strict $\lambda=0$ case since in pure SYK4 the ground state is not always unique. For $M=0\bmod 8$ (GOE), there is just one ground state with fixed fermion parity. In contrast, for GUE and GSE, $M\neq 0\bmod 8$, the presence of degeneracy allows for the superposition of the two ground states with opposite fermion parity, which can enhance the SRE. We numerically determine the maximum and minimum values of this linear combination  [^5] Given the two ground states $\rm GS_{1}$ and $\rm GS_{2}$ with opposite fermion parity and a generic linear combination |$\ket{\psi}=\sin\theta\ket{\rm GS_{1}}+e^{i\phi}\cos\theta\ket{\rm GS_{2}},$ (21) we numerically optimize $\langle\mathcal{M}_{2}(\psi)\rangle$ over a $200\times 200$ grid in the $(\theta,\phi)$ plane to determine the maximum and minimum values., as shown in the left inset of Fig. 6. Notably, the minima coincide with the values shown in the main plot (c) at small but non-zero $\lambda$, confirming that the Ising Hamiltonian breaks the degeneracy and selects one specific parity.
@@ VI Conclusion
The presence of distinct crossover values at finite system sizes is crucial for both classical and quantum simulations, as the feasible techniques are often constrained by available quantum resources. For instance, in noisy intermediate-scale quantum (NISQ) devices, the bottleneck is the ability to build up entanglement across the device Preskill (2012), while in tensor networks, the reachable bond dimension constrains the accessible entanglement entropy Eisert (2013); Orús (2014, 2019); Montangero et al. (2018). In contrast, the efficiency of classical calculations using the Gottesmann–Knill framework Gottesman (1997, 1998a); Aaronson and Gottesman (2004); Nest (2008); Gidney (2021) as well as the cost of broad classes of fault-tolerant quantum computers AI (2023); Katabarwa et al. (2024) is determined by the number of $T$-gates Bravyi et al. (2019); O’Gorman and Campbell (2017), which is related to the magic in a quantum state Li and Luo (2023). In this context, our analysis could guide the identification of the most efficient simulation strategies for each regime.
@@ Appendix A Mobility edge in the energy spectrum
[…]
@@ Appendix C Details on number of samples
Moreover, we adapted the number of samples for the marker under consideration. The computational complexity for computing the maximum entanglement grows as (${N\choose N/2}\times\text{construction}(\rho_{A})\times\text{diag}(\rho_{A})=\frac{2^{5/2N}}{\sqrt{N}}$. Similarly, the complexity for computing the Stabilizer Rényi Entropy (SRE) scales as $\mathcal{O}(2^{2N})$, due to the size of the Pauli group. In comparison, the computational complexity for calculating the fractal dimension is more modest, scaling as $\mathcal{O}(2^{N})$, permitting us to increase the number of samples to even further improve the quality of the results.
@@
Table 1: Number of samples used for fractal dimension estimates across different models and system sizes $N$.
@@
Table 2: Number of samples used for maximum entanglement entropy estimates across different models and system sizes $N$.
@@
Table 3: Number of samples used for Stabilizer Rényi Entropy estimates across different models and system sizes $N$.
@@ Appendix D Self-averaging property
[…]
@@
Systems are termed strongly self-averaging if $\beta\geq 1$, and weakly self-averaging when $10<\beta<$ 1. Notably, Fig. 8a2 shows that up to $\gamma=1.5$, already in the non-ergodic phase, the system remains strongly self-averaging. Beyond $\gamma=2.5$, all markers exhibit the same exponents. Additionally, $\mathcal{M}$ demonstrates stronger self-averaging in the ergodic regime and undergoes the most pronounced change around the transition points.
@@ Appendix F Pauli spectrum for RP and PLRBM bulk states
$\Pi_{\mathcal{P}}(\ket{\psi})=\sum_{i}\frac{\delta(x-x_{P_{i}})}{4^{N}}\,,\qquad x_{P_{i}}=\langle\psi|P_{i}|\psi\rangle,$ (30)
@@
Notably, exact stabilizer states yield Pauli expectation values of either $+1$ for all $2^{N}$ stabilizers, or a balanced set of $+1$ and $-1$ values Turkeshi et al. (2025a), as can be observed in the RP bulk case (blue lines in a2). In the PLRBM model, however, bulk localized states do not align with the computational basis $\{I,Z\}^{\otimes N}$, resulting in a more dispersed Pauli spectrum (b1) compared to the RP model (a1). Even after averaging over random realizations, the PLRBM bulk state retains high variance in its Pauli spectrum across the non-ergodic and localized regimes, except in the ergodic phase where the distribution sharply concentrates Vallejo-Fabila et al. (2024).
@@ Appendix G Complexity markers for superposition of two computational basis states
|$\ket{\psi}=\sin{\theta}\ket{a}+e^{i\phi}\cos{\theta}\ket{\bar{a}}\,,$ (31)
where |$\ket{a}$ and |$\ket{\bar{a}}$ are $N$-qubit orthogonal computational basis states differing by a Hamming distance of $N$, i.e., they are locally orthogonal at each qubit site. For such a state, the complexity markers are given by:
@@
Among the three markers, only the SRE is sensitive to the relative phase $\phi$. In the special case of an equal superposition with ±$\pm 1$ difference, i.e., $\sin{\theta}=\cos{\theta}=\frac{1}{\sqrt{2}}$ and $\phi=0,{\rm or\ }\pi$, the complexity markers simplify to
@@
|$\ket{\psi}=\sin{\theta}\ket{\rm GS_{1}}+e^{i\phi}\cos{\theta}\ket{\rm GS_{2}}\,.$ (37)
As |$\ket{\psi}$ should also be symmetric under $(-1)^{F}$ with possible eigenvalues ±$\pm 1$,
$(-1)^{F}\ket{\psi}\stackrel{{\scriptstyle!}}{{=}}\pm\ket{\psi}\Rightarrow\theta=\pi/4,\phi=0,\pi.$ (38)
@@ Appendix H Comparison between IPR and $D_{2}$
Although the fractal dimension $D_{2}$ can be formally related to the inverse participation ratio (IPR) $I_{2}$, defined in Eq. (1), via the expression $D_{q}=\frac{\log_{2}(I_{q})}{N(1-q)}$, relying solely on $I_{q}$ to detect phase transitions in such models can be challenging. This is because the multifractal nature of the wavefunctions manifests through the scaling relation $I_{q}\propto 1/L^{D_{q}(q-1)}$ Castellani and Peliti (1986); Mirlin et al. (1996); Mirlin and Evers (2000); Evers and Mirlin (2000); Čadež et al. (2024), which can be captured in $D_{q}$ instead of $I_{q}$. As illustrated in Fig. 11, the averaged IPR $\langle I_{2}\rangle$ reveals only the transition near $\gamma\simeq 2$, corresponding to the non-ergodic to localized transition, while it fails to resolve the ergodic to non-ergodic transition. This limitation highlights the advantage of using the fractal dimension as a more sensitive diagnostic of intermediate, multifractal phases.

### v2

@@ V.2 Results
In the integrable Ising model at $\lambda=1$, the ground state is doubly degenerate. However, introducing a small perturbation from the SYK Hamiltonian, $\lambda\lesssim 1$, lifts this degeneracy. Due to the fermion parity symmetry of the full Hamiltonian, the system selects a specific ground state: an equal superposition of the two antiferromagnetic spin configurations, with a relative phase of $\pm 1$. As a result, in this limit, the fractal dimension and maximal von-Neumann entropy density approach the values of $1/N$. In contrast, the magic tends to zero since this superposition is a stabilizer state (see Appendix G). All considered markers approach the values of the $\lambda\sim 1$ limit relatively rapidly without appreciable changes beyond $\lambda\gtrsim 0.5$, see Fig. 6.
@@
[…]
@@ VI Conclusion
The presence of distinct crossover values at finite system sizes is crucial for both classical and quantum simulations, as the feasible techniques are often constrained by available quantum resources. For instance, in noisy intermediate-scale quantum (NISQ) devices, the bottleneck is the ability to build up entanglement across the device Preskill (2012), while in tensor networks, the reachable bond dimension constrains the accessible entanglement entropy Eisert (2013); Orús (2014, 2019); Montangero et al. (2018). In contrast, the efficiency of classical calculations using the Gottesmann–Knill framework Gottesman (1997, 1998a); Aaronson and Gottesman (2004); Nest (2008); Gidney (2021) as well as the cost of broad classes of fault-tolerant quantum computers AI (2023); Katabarwa et al. (2024) is determined by the number of $T$-gates Bravyi et al. (2019); O’Gorman and Campbell (2017), which is related to the magic in a quantum state Howard and Campbell (2017). In this context, our analysis could guide the identification of the most efficient simulation strategies for each regime.
@@ Appendix A Mobility edge in the energy spectrum
The rapid convergence becomes even clearer when examining the evolution of the crossing points of the derivatives of the complexity markers, which we used in the main text to pinpoint the transitions. As shown in Fig. 7a2–c2, these crossing points quickly stabilize to a constant value above the energy window $10\%–15\%$, confirming the robustness of our choice.
@@ Appendix C Details on number of samples
Moreover, we adapted the number of samples for the marker under consideration. The computational complexity for computing the maximum entanglement grows as ${N\choose N/2}\times\text{construction}(\rho_{A})\times\text{diag}(\rho_{A})=\frac{2^{5/2N}}{\sqrt{N}}$. Similarly, the complexity for computing the Stabilizer Rényi Entropy (SRE) scales as $\mathcal{O}(2^{2N})$, due to the size of the Pauli group. In comparison, the computational complexity for calculating the fractal dimension is more modest, scaling as $\mathcal{O}(2^{N})$, permitting us to increase the number of samples to even further improve the quality of the results.
@@
Table 1: Number of samples used for fractal dimension estimates across different models and system sizes NN.
@@
Table 2: Number of samples used for maximum entanglement entropy estimates across different models and system sizes NN.
@@
Table 3: Number of samples used for Stabilizer Rényi Entropy estimates across different models and system sizes NN.
@@ Appendix D Self-averaging property
where $\langle\cdot\rangle$ denotes the disorder average. In ergodic phases, where the system uniformly explores its Hilbert space, $R_{\mathcal{C}}\to 0$ as the system size increases, indicating strong self-averaging. In contrast, in non-ergodic or localized regimes, $R_{\mathcal{C}}$ can remain finite even in the thermodynamic limit, reflecting strong sample-to-sample fluctuations and the breakdown of ergodicity.
@@
Systems are termed strongly self-averaging if $\beta\geq 1$, and weakly self-averaging when $0<\beta<1$. Notably, Fig. 8a2 shows that up to $\gamma=1.5$, already in the non-ergodic phase, the system remains strongly self-averaging. Beyond $\gamma=2.5$, all markers exhibit the same exponents. Additionally, $\mathcal{M}$ demonstrates stronger self-averaging in the ergodic regime and undergoes the most pronounced change around the transition points.
@@ Appendix F Pauli spectrum for RP and PLRBM bulk states
[…]
@@
Notably, exact stabilizer states yield Pauli expectation values of either +1+1 for all $2^{N}$ stabilizers, or a balanced set of +1+1 and 1-1 values Turkeshi et al. (2025a), as can be observed in the RP bulk case (blue lines in a2). In the PLRBM model, however, bulk localized states do not align with the computational basis $\{I,Z\}^{\otimes N}$, resulting in a more dispersed Pauli spectrum (b1) compared to the RP model (a1). Even after averaging over random realizations, the PLRBM bulk state retains high variance in its Pauli spectrum across the non-ergodic and localized regimes, except in the ergodic phase where the distribution sharply concentrates Vallejo-Fabila et al. (2024).
@@ Appendix G Complexity markers for superposition of two computational basis states
[…]
@@
Among the three markers, only the SRE is sensitive to the relative phase $\phi$. In the special case of an equal superposition with $\pm 1$ difference, i.e., $\sin{\theta}=\cos{\theta}=\frac{1}{\sqrt{2}}$ and $\phi=0,{\rm or\ }\pi$, the complexity markers simplify to
@@
[…]
@@ Appendix H Comparison between IPR and $D_{2}$
Although the fractal dimension $D_{2}$ can be formally related to the inverse participation ratio (IPR) $I_{2}$, defined in Eq. (1), via the expression $D_{q}=\frac{\log_{2}(I_{q})}{N(1-q)}$, relying solely on $I_{q}$ to detect phase transitions in such models can be challenging. This is because the multifractal nature of the wavefunctions manifests through the scaling relation $I_{q}\propto 1/L^{D_{q}(q-1)}$ Castellani and Peliti (1986); Mirlin et al. (1996); Mirlin and Evers (2000); Evers and Mirlin (2000); Čadež et al. (2024), which can be captured in $D_{q}$ instead of $I_{q}$. As illustrated in Fig. 11, the averaged IPR $\langle I_{2}\rangle$ reveals only the transition near $\gamma\simeq 2$, corresponding to the non-ergodic to localized transition, while it fails to resolve the ergodic to non-ergodic transition. This limitation highlights the advantage of using the fractal dimension as a more sensitive diagnostic of intermediate, multifractal phases. Moreover, taking the logarithm of the IPR is known to generally reduce the sample-to-sample fluctuations and improve the self-averaging behavior Solórzano et al. (2021).